OpenStudy (inexilewetrust):

Can someone Check my answer? What is the domain of the function f(x) = x2 + 5?

OpenStudy (inexilewetrust):

I chose the last one.

OpenStudy (inexilewetrust):

What is the range of the function f(x) = x2 + 5?

OpenStudy (inexilewetrust):

^^I chose the third one^^

OpenStudy (phi):

the domain are all the x values that are "legal" in the function. There is no number that "causes trouble" : e.g. 1) divide by 0, 2) sqrt of a negative number in other words, all real numbers are ok for the domain

OpenStudy (phi):

***What is the range of the function f(x) = x2 + 5?**** the range are all the numbers that f(x) can be notice that x^2 means x*x if x is negative: example x=-1, we would get -1*-1= 1 (a POSITIVE number) in other words, all numbers for x , after being squared , are positive. the smallest x^2 we can get is 0*0= 0 so the smallest f(x) can be is 0+5 = 5 Can you try again?

OpenStudy (inexilewetrust):

Okay, give me a few so I can think this over. THANK YOU SO MUCH

OpenStudy (inexilewetrust):

Quick question, what does the "R" mean?

OpenStudy (phi):

the script R is short for "all real numbers"

OpenStudy (inexilewetrust):

Okay so for the first one, it is the "R" one, correct?

OpenStudy (phi):

yes

OpenStudy (inexilewetrust):

I think the second one is real numbers, right?

OpenStudy (phi):

the range is not all real numbers. if you have x^2 + 5 how would you get , for example, 0 ? no matter what x you use, you won't get 0 out (for the reason I posted up above)

OpenStudy (inexilewetrust):

Okay, so the targeted number is zero?

OpenStudy (phi):

do you see how if x is 0, you get 0*0+ 5 or just 5 ? if x is bigger than 0 you get something bigger than 5 if x is less than 0, after you square it, it becomes positive, and you add this to 5... making the answer bigger than 5.

OpenStudy (phi):

**Okay, so the targeted number is zero? *** no. zero was just an example of a real number that you can't get, using the expression x*x + 5

OpenStudy (inexilewetrust):

I am not sure if zero or 5 is the best answer.

OpenStudy (inexilewetrust):

I am really sorry. I am really confused. Thank you for your patience.

OpenStudy (phi):

you are given f(x) = x*x+ 5 and are asked, what numbers can f(x) be? (i.e. what is the range of f(x) ) I am trying to show that the *smallest* f(x) can be is 5 (when x is 0) you can test a few x values: -1, 0, +1 you will get out -1 --> 6 0 --> 5 1 --> 6 and it should be clear that as x gets big, what you get out also gets big (any number we want... but we can never get smaller than 5)

OpenStudy (phi):

Yes, I can see you are missing some idea. If you can explain what is confusing, maybe we can figure out what idea you need to learn.

OpenStudy (inexilewetrust):

I am not sure what number I am trying to get for the question. What you explain seems easy but in a way not as 'straight' as I would like to put it.

OpenStudy (phi):

*** What is the range of the function f(x) = x2 + 5?*** The first idea is this: you are given a "rule" to "find numbers" The rule says: pick a number (in this case any real number) multiply it by itself add 5 the number we get is the "new number" and we call it f(x) the question *** What is the range*** is asking, "what numbers can f(x) be?" ok so far?

OpenStudy (inexilewetrust):

Yeah, so far so good!

OpenStudy (phi):

to find what numbers f(x) can be, we could do this: try all possible x values, and see what f(x) numbers we get out. that is a lot of work.... but we can use common sense. Let's think of just positive x numbers, from 0 to infinity as we make x bigger, the rule f(x) =x*x+5 gives us new numbers what number is f(x) when x is 0 ?

OpenStudy (phi):

do you know how to find f(x) when x is zero ?

OpenStudy (inexilewetrust):

No

OpenStudy (phi):

The rule says: pick a number (in this case we are doing zero) multiply it by itself add 5

OpenStudy (inexilewetrust):

Okay for -5 it would be 25+5=30

OpenStudy (phi):

yes. but let's just do x=0

OpenStudy (inexilewetrust):

Okay. so it would be five.

OpenStudy (phi):

yes. and for x=1 ?

OpenStudy (inexilewetrust):

6.

OpenStudy (phi):

and say we picked x= 0.5 (x= 1/2)

OpenStudy (phi):

0.5*0.5 = 0.25 (or 1/2 * 1/2 = 1/4 which is 0.25) then add +5 x= 1/2 gives us 5.25

OpenStudy (inexilewetrust):

Okay, so far so good!

OpenStudy (phi):

The idea is as x gets bigger, we can get *any* number out that is 5 or bigger does that make sense?

OpenStudy (inexilewetrust):

Yeah!

OpenStudy (phi):

now think about it. at x=0 we get out 5 for x bigger than 0, we do x*x (to get a number bigger than 0) and when we add 5, we get a number bigger than 5 so I expect that we can get numbers out from 5 up to whatever you want. what about if x is negative ?? i.e. x is less than zero?

OpenStudy (phi):

for example, what do we get for x= -1

OpenStudy (phi):

use the rule: -1 * -1 + 5 1+5 6 we get out 6 ok ?

OpenStudy (inexilewetrust):

Yeah, makes sense!

OpenStudy (phi):

if x is negative, and we multiply it by itself, we get a positive number (because a minus times a minus is plus) then we add 5 it looks like as x gets more and more negative, f(x) gets bigger and bigger ok?

OpenStudy (inexilewetrust):

Okay.

OpenStudy (phi):

so it looks like the "range" is 5 or bigger

OpenStudy (phi):

the allowed x numbers (i.e. we call it the domain) are all real numbers which means negative numbers, zero, and positive numbers. we figured out that for all these x's, the rule x*x + 5 can never be smaller than 5 but it can be as big as we want (by picking a big enough x)

OpenStudy (inexilewetrust):

Oh, wait, I think I hear the hamster turning the wheel in my head. I am sure the answer is D because it is adding positive five. So the X has to equal five or bigger...

OpenStudy (phi):

to write down the answer for the range we use \[ f(x) \ge 5 \] we put the f(x) next to the "big end" of > and the 5 next to the "small end" of >

OpenStudy (phi):

don't confuse x with f(x) x is what we put in (to the rule) f(x) is the new number we get out. we are figuring out what f(x) "ranges over"

OpenStudy (phi):

The reason we learn this stuff is for when we learn about inverse functions.

OpenStudy (inexilewetrust):

Oh.. got it.

OpenStudy (phi):

one of the reasons people are taught math is because it makes the hamster work hard. It needs exercise.

OpenStudy (inexilewetrust):

Yeah, mine is out of shape when it comes to math.